OpenStudy (anonymous):

Simplify completely √27y²

OpenStudy (saifoo.khan):

\[3 y \sqrt{3}\]

OpenStudy (rulnick):

3|y|sqrt(3). You need the absolute value around the y.

OpenStudy (anonymous):

\[\sqrt{9*3}*\sqrt{y^2}\] =\[3y \sqrt{3}\]

OpenStudy (anonymous):

i dont see the choice/

OpenStudy (mridul):

3\[3y \sqrt{3}\]

OpenStudy (anonymous):

and yeah there should be an absolute value sign!

OpenStudy (anonymous):

is it 3y power of 4√3

OpenStudy (rulnick):

If the problem is sqrt(27y^2) then the answer equals sqrt(27)sqrt(y^2) = sqrt(9*3)|y| = 3|y|sqrt(3).

OpenStudy (anonymous):

and sqrt(x^2) = |x| = +-x in other words if you graph sqrt(x^2) it equals a straight line in both the positive and negative directions (like an aboslute value graph)

OpenStudy (anonymous):

____ Simplify completely √27y²

OpenStudy (rulnick):

Yes, then the answer you seek is 3, times absolute value of y, times square root of 3.

OpenStudy (anonymous):

you guys are the best! :)

OpenStudy (anonymous):